OpenStudy (anonymous):

Find critical points and use interval notation to indicate where f(x) is increasing. For f(x)=(4+5x)/(3-2x)

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

What's the matter? Don't you want to know how to do it?

OpenStudy (anonymous):

f'(x) = (5(3-2x) - -2(4+5x))/(3 - 2x)^2 = 23/(3 - 2x)^2 a squared number is always non-negative, hence f'(x) > 0 for all x except 3/2 hence f is increasing for all real x except 3/2

OpenStudy (anonymous):

what about critical #'s?

OpenStudy (anonymous):

there's also a slanted asymptote somewhere between 1 and 2

OpenStudy (anonymous):

Okay, yeah, jamesm is right

OpenStudy (anonymous):

critical point: take the derivative find out where it is 0 find out where it is undefined. those are the critical points.

OpenStudy (anonymous):

derivative is \[\frac{23}{(3-2x)^2}\]

OpenStudy (anonymous):

this is never zero because the numerator is a constant. it is undefined at \[x=\frac{3}{2}\] but then again so is your original function so you get no information from this

OpenStudy (anonymous):

increasing where the derivative is positive, and this is ALWAYS positive since numerator is a positive constant and the denominator is a square

OpenStudy (anonymous):

there is no slant asympote so ignore that. the degree of the numerator is equal the degree of the denominator so you have a horizontal asympote at \[y=-\frac{5}{2}\] the ratio of the leading coefficients

OpenStudy (anonymous):

vertical asymptote at \[x=\frac{3}{2}\]

OpenStudy (anonymous):

so if you draw those two asympotes and remember that your function is always increasing you only have one picture you can draw.

OpenStudy (anonymous):

satellite...it is always increasing where f is defined, yes? can't say it's always increasing in general